OpenStudy (anonymous):

Write a compound inequality for each problem: 11. The human ear can distinguish sounds between 20 Hz and 20,000 HZ; inclusive 12. For a man to box as a welterweight, he must weigh more then 140 LBS, but at most 147 LBS.

OpenStudy (anonymous):

\[20 \le x \le 20000\] \[140 < x \le 147\]

OpenStudy (kainui):

Just consider what you're looking at. Is it greater than or equal to? Is it less than but not equal to?